# Berms without Calibration

### v1

@@ Abstract
We derive a new semi-analytical pricing model for Bermudan swaptions based on swap rates distributions and correlations between them. The model does not require product specific calibration.
@@ Introduction
A Bermudan swaption on an interest rate swap gives the holder the right to enter the swap on some specified exercise dates. This product is considered to be the simplest exotic which requires the use of the term structure models for valuations and risk management. It is customary to use an LGM style model to price Bermudan swaptions. The volatility curve of the underlying Hull-White process is calibrated to the implied volatilities of the coterminal swaptions and the mean reversion curve is often recalibrated to the results of the Totem submissions. It is known that the mean reversion parameter is related to the forward volatility of the swaptions par rates (see for example [2]) but calibration techniques with forward volatilities are not widely used, as it is difficult to express the forward volatility in terms of products liquidly traded in the interest rate market. Often midcurve swaptions (swaptions with a significant delay between the swaption expiry and the underlying start date) are considered as products on the forward volatility of the underlying (forward starting) swap rate, though fitting them in the term structure model calibration routine is not an easy task.
@@
In this paper we connect valuations of Bermudan swaptions and midcurve swaptions directly. As a connecting bridge we look at a less known modification of the midcurve swaption - the swaption with a relative strike where the strike is fixed sometime before the swaption exercise, but the swaption exercise is standard and within few business days of the underlying start. This swaption contract directly depends on the forward volatility between the strike fixing date and the swaption exercise date. To clarify the midcurve-Bermudan swaptions relationship we study first the Canary swaption [9] which is the simplest case of a Bermudan swaption - the swaption with only two exercise dates. Representing a Canary swaption as a basket of two option style products on coterminal swap rates we derive an analytic valuation based on minor assumptions on the dynamics of the swap rates between the first and the second expiry. We generalise the valuation to the case of arbitrary Bermudan swaption by developing Hagan LGM [7] style model.
@@
The main result of the present paper is similar in nature to the classical reduction of an European swaption valuation to the product of the swap annuity and a vanilla option price. We show how the numeraire change from the zero coupon bond to the annuity simplifies valuations of Bermudan swaptions by replacing unobservable mean reversion parameters with swap rates correlations. The later can be easily estimated from the historical data, or directly implied from midcurve swaptions or CMS spreads markets.
@@
The model developed in the present paper can be used to improve risk management of Bermudans as in addition to vanilla swaptions it allows inclusion of the simplest IR correlation products in to the hedging portfolio. Another advantage of the approach is that it does not require any preliminary product specific calibration. We also provide a toy model under the assumption of deterministic ratios of annuities and perfect correlations of the forward swap rates. In this setting we show how to value a Bermudan swaption as a zero floor on the maximum of correlated Gaussian variables using techniques presented in [10].
@@
The paper has the following structure. In Section 1 we introduce the notations to be used throughout the paper, explain the pricing mechanism for midcurve swaptions and formulate the modelling assumption on annuity ratios. In Section 2 we introduce a modification to the swaption contract - swaptions with relative strikes, and provide its valuation formula. In Section 3 we present explicit valuation formulas for Canary swaptions under the assumption of deterministic ratios of annuities. In Section 4 we discuss an extension of the Canary swaption pricing approach to the case of a generic Bermudan swaption. We provide a general valuation mechanism and give two practical specialisations - one suitable for the moment matching valuation and another for Hagan lattice valuation. We discuss how stochastic annuity ratios can be handled in Section 5. We summarise our findings in Conclusion.
@@ 1 Midcurve Swaptions
This section follows closely [5] and we refer the reader to it for all the main details, notations and proofs. We shall be working in the annuity probability measure $\cal A$} where the relevant underlying swap annuity is a martingale. We assume that the corresponding sample space of continuous outcomes has a time dependent filtration $\{{\cal F}_{s},s\in[t,T_{e}]\}$ with ${\cal F}_{s_{1}}\subseteq{\cal F}_{s_{2}}$, whenever $s_{1}\leq s_{2}$, $t$ being the time corresponding to the valuation date and $T_{e}$ corresponding to the end date of the underlying swap (see [1] for details of the use of martingale theory for interest rate modelling).
A midcurve swaption is a swaption $W(t,T_{ex},S(T_{s},T_{e},K))$ on a forward starting swap $S(T_{s},T_{e},K)$ with a significant delay between the swaption expiry $T_{ex}$ and the underlying swap start date $T_{s}$ ($T_{ex}<T_{s}$, $K$ is the swaption strike which coincides with the underlying swap fixed rate). The underlying (forward starting) swap $S(T_{s},T_{e},K)$ can be decomposed in to the difference of the long swap $S(T_{ex},T_{e},K)$ from the expiry date $T_{ex}$ to the underlying swap end date $T_{e}$ and the short swap $S(T_{ex},T_{s},K)$ from the expiry date $T_{ex}$ to the underlying swap start date $T_{s}$. This decomposition allows a midcurve swaption to be priced in the underlying swap annuity measure as an option on a correlated weighted basket of the long swap par rate $R(t,T_{ex},T_{e})$ and the short swap par rate $R(t,T_{ex},T_{s})$ with the weights given by the ratios of the respective swap annuities - the long swap annuity $A(t,T_{ex},T_{e})$, the short swap annuity $A(t,T_{ex},T_{s})$ and the underlying swap annuity $A(t,T_{s},T_{e})$ (see [5] for more detailed explanation):
@@
$=$ $A(t,T_{s},T_{e})\mathbb{E}^{\cal A}\Bigr[\Bigl(\omega\Bigl(\frac{A(T_{ex},T_{ex},T_{e})}{A(T_{ex},T_{s},T_{e})}R(T_{ex},T_{ex},T_{e})$-
$\frac{A(T_{ex},T_{ex},T_{s})}{A(T_{ex},T_{s},T_{e})}R(T_{ex},T_{ex},T_{s})-K\Bigr)\Bigr)^{+}\Big|{\cal F}_{t}\Bigr].$
Here and in what follows $\omega=\pm 1$ is used to denote whether the option is a payer or a receiver. We shall use the first argument in $A(\cdot,T_{s},T_{e})$ and $R(\cdot,T_{s},T_{e})$ to specify the observation (or the fixing) time of the stochastic variable. We reserve the low case $t$ to specify the valuation date so that $A(t,T_{s},T_{e})$ and $R(t,T_{s},T_{e})$ are the expected (fair) values of the swap annuity and the swap par rate at the valuation date $t$. The upper case $T$ will be used to indicate that the variables $A(T,T_{s},T_{e})$, $R(T,T_{s},T_{e})$ are stochastic when seen from the valuation date $t$. As both $A(T,T_{s},T_{e})$, $R(T,T_{s},T_{e})$ are martingales in the underlying (forward starting) swap annuity measure we can write:
@@
$\tilde{x}$ $=$ $\frac{A(t,T_{s},T_{e})}{A(t,T_{ex},T_{e})}\mathbb{E}^{\cal A}\left[\frac{A(T_{ex},T_{ex},T_{e})}{A(T_{ex},T_{s},T_{e})}\Big|x=R(T_{ex},T_{ex},T_{e}),y=R(T_{ex},T_{ex},T_{s}),{\cal F}_{T_{ex}}\right]x=f(x,y)x,$
$\tilde{y}$ $=$ $\frac{A(t,T_{s},T_{e})}{A(t,T_{ex},T_{s})}\mathbb{E}^{\cal A}\left[\frac{A(T_{ex},T_{ex},T_{s})}{A(T_{ex},T_{s},T_{e})}\Big|x=R(T_{ex},T_{ex},T_{e}),y=R(T_{ex},T_{ex},T_{s}),{\cal F}_{T_{ex}}\right]y=g(x,y)y,$
@@
$W(t,T_{ex},S(T_{s},T_{e},K))$ $=$ $A(t,T_{s},T_{e})\int^{\infty}_{-\infty}\int^{\infty}_{-\infty}(a\tilde{x}-b\tilde{y}-K)^{+}\Psi(\Psi^{u}_{e}(x),\Psi^{u}_{s}(y))pdf^{u}_{e}(x)pdf^{u}_{s}(y)\times$ (7)
$\times$ $det\left(\begin{array}[]{cc}\frac{\partial\tilde{x}}{\partial x}&\frac{\partial\tilde{y}}{\partial x}\\ \frac{\partial\tilde{x}}{\partial y}&\frac{\partial\tilde{y}}{\partial y}\end{array}\right)^{-1}d\tilde{x}d\tilde{y}$,
@@
where $\Psi$ is the copula kernel and $\Psi^{u}_{e},\Psi^{u}_{s}$ are the cumulative density functions of the swap rates distributions $pdf^{u}_{e},pdf^{u}_{s}$ in the underlying (forward starting) swap annuity measures. The integration weight can be simplified further
$det\left(\begin{array}[]{cc}\frac{\partial\tilde{x}}{\partial x}&\frac{\partial\tilde{y}}{\partial x}\\ \frac{\partial\tilde{x}}{\partial y}&\frac{\partial\tilde{y}}{\partial y}\end{array}\right)^{-1}$ $=$ $det\left(\begin{array}[]{cc}f+xf_{x}&yg_{x}\\ xf_{y}&g+yg_{y}\end{array}\right)^{-1}$
$=$ $f(\cdot)^{-1}g(\cdot)^{-1}\left(1+x\frac{f_{x}}{f}+y\frac{g_{y}}{g}+xy\frac{f_{x}g_{y}-f_{y}g_{x}}{fg}\right)^{-1}$
$=$ $f(\cdot)^{-1}g(\cdot)^{-1}\left(1+\tilde{x}\frac{f_{x}}{f^{2}}+\tilde{y}\frac{g_{y}}{g^{2}}+\tilde{x}\tilde{y}\frac{f_{x}g_{y}-f_{y}g_{x}}{f^{2}g^{2}}\right)^{-1}.$ (13)
If we assume that annuity ratios $\frac{A(T,T_{ex},T_{e})}{A(T,T_{s},T_{e})}$ and $\frac{A(T,T_{ex},T_{s})}{A(T,T_{s},T_{e})}$ are deterministic then the weight factor $fg\left(1+\tilde{x}\frac{f_{x}}{f^{2}}+\tilde{y}\frac{g_{y}}{g^{2}}+\tilde{x}\tilde{y}\frac{f_{x}g_{y}-f_{y}g_{x}}{f^{2}g^{2}}\right)$ is 1. In the general case it is convenient to linearise the weight in such a way that the following assumption holds
@@
$\Psi(\Psi^{u}_{e}(x),\Psi^{u}_{s}(y))f(\cdot)^{-1}g(\cdot)^{-1}\left(1+\tilde{x}\frac{f_{x}}{f^{2}}+\tilde{y}\frac{g_{y}}{g^{2}}+\tilde{x}\tilde{y}\frac{f_{x}g_{y}-f_{y}g_{x}}{f^{2}g^{2}}\right)^{-1}$ $pdf^{u}_{e}(x)pdf^{u}_{s}(y)$ $\approx$
@@
Practitioners often manage midcurve swaptions in terms of midcurve implied volatility-by-strike which is approximated as
$\sigma^{2}_{s,e}(K)$ $=$ $\frac{A(t,T_{ex},T_{e})^{2}}{A(t,T_{s},T_{e})^{2}}\sigma^{2}_{e}(K_{e})-2\frac{A(t,T_{ex},T_{e})}{A(t,T_{s},T_{e})}\frac{A(t,T_{ex},T_{s})}{A(t,T_{s},T_{e})}\rho_{s,e}\sigma_{e}(K_{e})\sigma_{s}(K_{s})+\frac{A(t,T_{ex},T_{s})^{2}}{A(t,T_{s},T_{e})^{2}}\sigma^{2}_{s}(K_{s}),$ (16)
$K_{e}=R(t,T_{ex},T_{e})+(K-R(t,T_{s},T_{e})),$
$K_{s}=R(t,T_{ex},T_{s})+(K-R(t,T_{s},T_{e})),$
where $\rho_{s,e}$ is the implied correlation of the long and short swap rates as fixed at $T_{ex}$. We shall use this approximation (omitting the strike $K$ from the notations) to present results in the commonly used terminology of implied volatilities, while the procedure can be made rigorous by switching to local volatilities and using Andreasen-Huge one-time-step method [3], or using Black/SABR baskets approach from [4, 6, 8].
In the following sections we will rely on the following consequences of [5] and linearisation (LABEL:LNZ):
@@
Under the assumption of deterministic ratios of annuities, the Radom-Nikodym derivative of the measure changes when moving between the long, the short and the forward starting annuity measures are all 1. Thus, we can use distributions of the corresponding swap rates as observed in their natural measures and substitute them to the joint distribution of those rates in each of the measure.
Proof: The result follows from Lemma 1 [5] and formula (20) from Deterministic Annuity Ratio subsection of Section 2 in [5].
@@
Under the linearisation assumption (LABEL:LNZ) we can make convexity adjustments to the relevant swap rates so that the stochastic annuity ratio can be assumed deterministic.
Proof: This is a direct consequence of the integration formula (7), linearisation assumption (LABEL:LNZ) and Radon-Nikodym measure change.

### v2

@@ Abstract
We propose a new semi-analytical pricing model for Bermudan swaptions based on swap rate distributions and correlations between them. The model does not require product specific calibration.
@@ Introduction
A Bermudan swaption on an interest rate swap gives the holder the right to enter the swap on some specified exercise dates. This product is considered to be the simplest exotic which requires the use of a term structure model for valuations and risk management. It is customary to use an LGM style model to price Bermudan swaptions. The volatility curve of the underlying Hull-White process is calibrated to the implied volatilities of the coterminal swaptions; and the mean reversion curve is often recalibrated to the results of the Totem submissions. It is known that the mean reversion parameter is related to the forward volatility of the swaption par rates (see for example [2]), but calibration techniques with forward volatilities are not widely used, as it is difficult to express the forward volatility in terms of products that are liquidly traded in the interest rate market. Often midcurve swaptions (swaptions with a significant delay between the swaption expiry date and the underlying start date) are considered as products on the forward volatility of the underlying (forward starting) swap rate, though fitting them in the term structure model calibration routine is not an easy task.
@@
In this paper we connect valuations of Bermudan swaptions and midcurve swaptions directly. As a connecting bridge, we look at a lesser known modification of the midcurve swaption - the swaption with a relative strike where the strike is fixed sometime before the swaption exercise, but the swaption exercise is standard and within few business days of the underlying start. This swaption contract directly depends on the forward volatility between the strike fixing date and the swaption exercise date. To clarify the midcurve-Bermudan swaption relationship we study first the Canary swaption [10], which is the simplest case of a Bermudan swaption - the swaption with only two exercise dates. Representing a Canary swaption as a basket of two option style products on coterminal swap rates, we derive an analytic valuation based on minor assumptions on the dynamics of the swap rates between the first and the second expiry. We generalise the valuation to the case of arbitrary Bermudan swaption by developing a Hagan LGM [8] style model.
@@
The main result of the present paper is similar in nature to the classical reduction of a European swaption valuation to the product of the swap annuity and a vanilla option price. We show how the numeraire change from the zero coupon bond to the annuity simplifies valuations of Bermudan swaptions by replacing mean reversion parameters with swap rate correlations. The latter can be easily estimated from historical data, or directly implied from midcurve swaption or CMS spread markets.
@@
The model developed in the present paper can be used to improve risk management of Bermudans, as in addition to vanilla swaptions, it allows inclusion of the simplest IR correlation products into the hedging portfolio. Another advantage of the approach is that it does not require any preliminary product specific calibration. We also provide a toy model under the assumption of deterministic ratios of annuities and perfect correlations of the forward swap rates. In this setting we show how to value a Bermudan swaption as a zero floor on the maximum of correlated Gaussian variables using techniques presented in [11].
@@
The paper has the following structure. In Section 1, we introduce the notations to be used throughout the paper, explain the pricing mechanism for midcurve swaptions and formulate the modelling assumption on annuity ratios. In Section 2, we introduce a modification to the swaption contract - swaptions with relative strikes, and provide its valuation formula. In Section 3, we present explicit valuation formulas for Canary swaptions under the assumption of deterministic ratios of annuities. In Section 4, we discuss an extension of the Canary swaption pricing approach to the case of a generic Bermudan swaption. We provide a general valuation mechanism and give two practical specialisations - one suitable for the moment matching valuation, and another for Hagan lattice valuation. We discuss how stochastic annuity ratios can be handled in Section 5. We summarise our findings in Conclusion.
@@ 1 Midcurve Swaptions
This section follows [6] closely, and we refer the reader to it for the main details, notations and proofs. We will be working in the annuity probability measure $\cal A$} where the relevant underlying swap annuity is a martingale. We assume that the corresponding sample space of continuous outcomes has a time dependent filtration $\{{\cal F}_{s},s\in[t,T_{e}]\}$ with ${\cal F}_{s_{1}}\subseteq{\cal F}_{s_{2}}$, whenever $s_{1}\leq s_{2}$, $t$ being the time corresponding to the valuation date and $T_{e}$ corresponding to the end date of the underlying swap (see [1] for details of the use of martingale theory for interest rate modelling).
A midcurve swaption is a swaption $W(t,T_{ex},S(T_{s},T_{e},K))$ on a forward starting swap $S(T_{s},T_{e},K)$ with a significant delay between the swaption expiry $T_{ex}$ and the underlying swap start date $T_{s}$ ($T_{ex}<T_{s}$, $K$ is the swaption strike which coincides with the underlying swap fixed rate). The underlying (forward starting) swap $S(T_{s},T_{e},K)$ can be decomposed in to the difference of the long swap $S(T_{ex},T_{e},K)$ from the expiry date $T_{ex}$ to the underlying swap end date $T_{e}$ and the short swap $S(T_{ex},T_{s},K)$ from the expiry date $T_{ex}$ to the underlying swap start date $T_{s}$. This decomposition allows a midcurve swaption to be priced in the underlying swap annuity measure as an option on a correlated weighted basket of the long swap par rate $R(T,T_{ex},T_{e})$ and the short swap par rate $R(T,T_{ex},T_{s})$ with the weights given by the ratios of the respective swap annuities - the long swap annuity $A(T,T_{ex},T_{e})$ and the short swap annuity $A(T,T_{ex},T_{s})$ to the underlying swap annuity $A(T,T_{s},T_{e})$ (see [6] for more detailed explanation):
@@
$=$ $A(t,T_{s},T_{e})\mathbb{E}^{\cal A}\Bigr[\Bigl(\omega\Bigl(\frac{A(T_{ex},T_{ex},T_{e})}{A(T_{ex},T_{s},T_{e})}\Bigl(R(T_{ex},T_{ex},T_{e})-K\Bigr)$-
$\frac{A(T_{ex},T_{ex},T_{s})}{A(T_{ex},T_{s},T_{e})}\Bigl(R(T_{ex},T_{ex},T_{s})-K\Bigr)\Bigr)^{+}\Big|{\cal F}_{t}\Bigr].$
Here and in what follows $\omega=\pm 1$ is used to denote whether the option is a payer or a receiver. We shall use the first argument in $A(\cdot,T_{s},T_{e})$ and $R(\cdot,T_{s},T_{e})$ to specify the observation (or the fixing) time of the stochastic variable. We reserve the low case $t$ to specify the valuation date so that $A(t,T_{s},T_{e})$ and $R(t,T_{s},T_{e})$ are the expected (fair) values of the swap annuity and the swap par rate at the valuation date $t$. The upper case $T$ will be used to indicate that the variables $A(T,T_{s},T_{e})$, $R(T,T_{s},T_{e})$ are stochastic when seen from the valuation date $t$. As both $A(T,T_{s},T_{e})$, $R(T,T_{s},T_{e})$ are martingales in the underlying (forward starting) swap annuity measure $\cal A$ we can write:
@@
$\tilde{x}$ $=$ $(x-K)\frac{A(t,T_{s},T_{e})}{A(t,T_{ex},T_{e})}\mathbb{E}^{\cal A}\left[\frac{A(T_{ex},T_{ex},T_{e})}{A(T_{ex},T_{s},T_{e})}\Big|x=R(T_{ex},T_{ex},T_{e}),y=R(T_{ex},T_{ex},T_{s}),{\cal F}_{T_{ex}}\right]$
$=$ $(x-K)f(x,y),$
$\tilde{y}$ $=$ $(y-K)\frac{A(t,T_{s},T_{e})}{A(t,T_{ex},T_{s})}\mathbb{E}^{\cal A}\left[\frac{A(T_{ex},T_{ex},T_{s})}{A(T_{ex},T_{s},T_{e})}\Big|x=R(T_{ex},T_{ex},T_{e}),y=R(T_{ex},T_{ex},T_{s}),{\cal F}_{T_{ex}}\right]$
$=$ $(y-K)g(x,y),$
@@
$W(t,T_{ex},S(T_{s},T_{e},K))$ $=$ $A(t,T_{s},T_{e})\int^{\infty}_{-\infty}\int^{\infty}_{-\infty}(a\tilde{x}-b\tilde{y})^{+}\Psi(\Psi^{u}_{e}(x),\Psi^{u}_{s}(y))pdf^{u}_{e}(x)pdf^{u}_{s}(y)\times$ (7)
$det\left(\begin{array}[]{cc}\frac{\partial\tilde{x}}{\partial x}&\frac{\partial\tilde{y}}{\partial x}\\ \frac{\partial\tilde{x}}{\partial y}&\frac{\partial\tilde{y}}{\partial y}\end{array}\right)^{-1}d\tilde{x}d\tilde{y}$,
@@
where $\Psi$ is the copula kernel and $\Psi^{u}_{e},\Psi^{u}_{s}$ are the cumulative density functions of the swap rates distributions $pdf^{u}_{e},pdf^{u}_{s}$ in the underlying (forward starting) swap annuity measures. The integration weight can be simplified further:
$det\left(\begin{array}[]{cc}\frac{\partial\tilde{x}}{\partial x}&\frac{\partial\tilde{y}}{\partial x}\\ \frac{\partial\tilde{x}}{\partial y}&\frac{\partial\tilde{y}}{\partial y}\end{array}\right)^{-1}$ $=$ $det\left(\begin{array}[]{cc}f+(x-K)f_{x}&(y-K)g_{x}\\ (x-K)f_{y}&g+(y-K)g_{y}\end{array}\right)^{-1}$
$=$ $f(\cdot)^{-1}g(\cdot)^{-1}\left(1+(x-K)\frac{f_{x}}{f}+(y-K)\frac{g_{y}}{g}+(x-K)(y-K)\frac{f_{x}g_{y}-f_{y}g_{x}}{fg}\right)^{-1}$
$=$ $f(\cdot)^{-1}g(\cdot)^{-1}\left(1+(x-K)\frac{f_{x}}{f}+(y-K)\frac{g_{y}}{g}\right)^{-1},$ (13)
where for the last transformation we used that $af(x,y)=bg(x,y)+1$, which is a consequence of the following identity:
$\frac{A(T_{ex},T_{ex},T_{e})}{A(T_{ex},T_{s},T_{e})}-\frac{A(T_{ex},T_{ex},T_{s})}{A(T_{ex},T_{s},T_{e})}$ $\equiv$ $1.$ (14)
If we assume that annuity ratios $\frac{A(T,T_{ex},T_{e})}{A(T,T_{s},T_{e})}$ and $\frac{A(T,T_{ex},T_{s})}{A(T,T_{s},T_{e})}$ are deterministic, then the weight factor $fg\left(1+(x-K)\frac{f_{x}}{f}+(y-K)\frac{g_{y}}{g}\right)$ is 1. In the general case, it is convenient to linearise the weight in such a way that the following assumption holds:
@@
We assume that there exist such deformations of the probability densities $pdf^{u}_{e}(x)$ and $pdf^{u}_{s}(y)$ that
@@
In the following sections we will rely on the following consequences of [6] and linearisation (15):
@@
Under the assumption of deterministic ratios of annuities, the Radom-Nikodym derivative of the measure changes when moving between the long, the short and the forward starting swap annuity measures are all equal to 1. Thus, we can use distributions of the corresponding swap rates as observed in their natural measures and substitute them to the joint distribution of those rates in each of the measure.
Proof: Using Lemma 1 from [6] we can relate the probability densities $pdf_{e}(x)$, $pdf_{s}(y)$ of the long and the short swap rates in their natural annuities measures to the probability densities $pdf^{u}_{e}(x)$, $pdf^{u}_{s}(y)$ of the same swap rates in the underlying (forward starting) swap annuity measure as:
$pdf^{u}_{e}(x)$ $=$ $pdf_{e}(x)\frac{A(t,T_{ex},T_{e})}{A(t,T_{s},T_{e})}\mathbb{E}^{{\cal A}_{e}}\left[\frac{A(T_{ex},T_{s},T_{e})}{A(T_{ex},T_{ex},T_{e})}\Big|x=R(T_{ex},T_{ex},T_{e}),{\cal F}_{T_{ex}}\right],$
$pdf^{u}_{s}(y)$ $=$ $pdf_{s}(y)\frac{A(t,T_{ex},T_{s})}{A(t,T_{s},T_{e})}\mathbb{E}^{{\cal A}_{s}}\left[\frac{A(T_{ex},T_{s},T_{e})}{A(T_{ex},T_{ex},T_{s})}\Big|y=R(T_{ex},T_{ex},T_{s}),{\cal F}_{T_{ex}}\right],$ (17)
where ${\cal A}_{e}$ and ${\cal A}_{s}$ are the long and the short swap annuity measures respectively. If the annuity ratios are deterministic then the terms under expectations cancel the normalising coefficients in front of them. This gives the result of the theorem.
@@
Under the linearisation assumption (15) we can make convexity adjustments to the relevant swap rates so that payoffs expressed as functions of
$\frac{A(T_{ex},T_{ex},T_{e})}{A(T_{ex},T_{s},T_{e})}(R(T_{ex},T_{ex},T_{e})-K)$ ${\rm and}$ $\frac{A(T_{ex},T_{ex},T_{s})}{A(T_{ex},T_{s},T_{e})}(R(T_{ex},T_{ex},T_{s})-K)$ (18)
can be valued using the deterministic annuity ratio assumption.
[…]
Practitioners often manage midcurve swaptions in terms of the midcurve implied volatility-by-strike which is approximated as
$\sigma^{2}_{s,e}(K)$ $=$ $\frac{A(t,T_{ex},T_{e})^{2}}{A(t,T_{s},T_{e})^{2}}\sigma^{2}_{e}(K_{e})-2\frac{A(t,T_{ex},T_{e})}{A(t,T_{s},T_{e})}\frac{A(t,T_{ex},T_{s})}{A(t,T_{s},T_{e})}\rho_{s,e}\sigma_{e}(K_{e})\sigma_{s}(K_{s})+\frac{A(t,T_{ex},T_{s})^{2}}{A(t,T_{s},T_{e})^{2}}\sigma^{2}_{s}(K_{s}),$
where $K_{e}=R(t,T_{ex},T_{e})+(K-R(t,T_{s},T_{e}))$, $K_{s}=R(t,T_{ex},T_{s})+(K-R(t,T_{s},T_{e}))$ and $\rho_{s,e}$ is the implied correlation of the long and short swap rates as fixed at $T_{ex}$. We shall use this approximation (omitting the strike $K$ from the notations) to present results in the commonly used terminology of implied volatilities, while the procedure can be made rigorous by switching to local volatilities and using Andreasen-Huge one-time-step method [4], or using Black/SABR baskets approach from [5, 7, 9].